A train approaching a station does two successive half-kilometers in 16 and 20 seconds respectively. Assuming the retardation to be uniform, find the further distance the train runs before stopping.
1
Expert's answer
2019-03-19T12:52:43-0400

We're going to use the fact that "s_1 = s_2 = 0.5~\text{km} = s" to simplify some calculations.

Given times: "t_1 = 16~\text{s}, t_2 = 20~\text{s}."

Let's assume the initial position of the train as "s_0 = 0." Then, the kinematic equations for the given reference positions:

"s = v_0t_1 + \frac{a}{2}t_1^2; \\
2s = v_0(t_1+t_2) + \frac{a}{2}(t_1+t_2)^2; \\
s_{stop} = v_0t_{stop} + \frac{a}{2}t_{stop}^2; \\
v_{stop} = 0 = v_0t_{stop} + at_{stop}."

Please pay attention that here the auxiliary values "s_{stop}, t_{stop}" are counted from the initial position and time.

By excluding the values from the above system of equations,

"v_0 = s\frac{t_2^2+2t_1t_2-t_1^2}{t_1t_2(t_1+t_2)}; \\
a = -2s\frac{t_2-t_1}{t_2(t_2^2+3t_1t_2+2t_1^2)}; \\
t_{stop} = -\frac{v_0}{a}; \\
s_{stop} = -\frac{v_0^2}{2a} = \frac{1}{4}s\frac{(t_2^2+3t_1t_2+2t_1^2)(t_2^2+2t_1t_2-t_1^2)^2}{t_1^2t_2(t_2-t_1)(t_1+t_2)^2}."

One can see that a is negative, which indicates retardation.

The further distance (after the first two portions) the train passed will be given by

"{\Delta}s = s_{stop} - 2s. \\"

By entering the numerical values into the formula, and canceling all the seconds units,

"{\Delta}s = 0.5~\text{km}~*~(\frac{1}{4}\frac{(20^2+3*16*20+2*16^2)(20^2+2*16*20-16^2)^2}{16^220(20-16)(16+20)^2} - 2) \approx 4.42~\text{km}."